12 June 2010

Saturday night with French cuisine

The inner part of the chip chamber pulled out of the drill for cleaning
The inner part of the chip chamber pulled out of the drill for cleaning. The metal rod is our hollow shaft. Attached to the hollow shaft are two soggy looking plastic bags. These bags contained the fluid with the beads. After deployment of the drill to the bottom of
the hole and
subsequent starting of a drilling run, the bags were ripped open, allowing the beads to be dispersed in the fluid while drilling. In the back ground: Chinese driller, Wang.

Weather continues to be really poor. Today we had overcast all day, very warm temperature and little wind. Snow is sticky and on all non-white surfaces pools of water form. It appears, that the weather pattern from last winter with above normal temperatures in Greenland and below normal temperatures in Eastern North America
and in Northern Europe continues. In Kangerlussuaq, Greenland, it is now warmer than in Copenhagen, Denmark.

At NEEM camp work in the trenches continues independent of weather. We are still able to maintain cold temperatures in the snow caves; but we would like some colder temperature soon.

It was the first Saturday evening since last crew exchange and dinner was prepared by a French/Belgian/U.S. team (Romain, Gregory, Sonia, Denis and Lou).

What we have done today:
1. Drilling and logging.
2. Processed ice cores: 26 bags, from 3584 to 3609.
3. Measuring CFA. Measured 6.6 m. CFA depth: 1612.05 m.
4. Grooming skiway with beam groomer.
5. Removing snow drifts between the two workshop tents.
6. Successful deployment of beads in the drilling fluid at the bottom of the bore hole.
    Subsequent ice cores will now be analyzed for how far contamination from drill fluid
    penetrates into the ice core.
7. Canadian and Japanese pit studies for biological material in snow.

Ad.1: Drillers report:
We had 3 runs today, including the run with bead deployment.

Driller's depth 2040.26 meters. Logging depth: 2053.15 m.

Ad.3: The production is a bit less today as a breakdown in the fast IC instrument caused the night shift to stop. The fast IC has been repaired and production is resumed.

Ad.6: Today we launched micro-florescent beads during drill run 707 to monitor infiltration into the core as part of a microbiological contamination study. All totaled, some 100 billion beads (1 micron) were introduced to the bottom of the hole using the
"cognac bomb" delivery technique. The bomb was made into three sausage shaped bags that were inserted into the chip chamber and released using the "Hansen Hook" developed last year. All three bags were successfully ripped and all three hooks recovered along with 3.3 m of core. Sections of the core will be returned for infiltration studies using a microscope and GC detection of drill fluid contamination.

Weather: Overcast with varying thickness. Temp. -7 °C to -3 °C, 5-10 knots from SE and SW. Visibility: Mostly unrestricted.

FL, J.P. Steffensen


← Previous entry   Next entry →